Following are several figures reported for Allister and Barone as of December 31, 2018:

Allister Barone
Inventory $ 490,000 $ 290,000
Sales 980,000 780,000
Investment income not given
Cost of goods sold 490,000 390,000
Operating expenses 225,000 295,000

Allister acquired 80 percent of Barone in January 2017. In allocating the newly acquired subsidiary's fair value at the acquisition date, Allister noted that Barone had developed a customer list worth $76,000 that was unrecorded on its accounting records and had a 5-year remaining life. Any remaining excess fair value over Barone's book value was attributed to goodwill. During 2018, Barone sells inventory costing $129,000 to Allister for $178,000. Of this amount, 20 percent remains unsold in Allister's warehouse at year-end.

Determine balances for the following items that would appear on Allister's consolidated financial statements for 2018:
Answer #1

Answer:

Particulars Amount Working notes
INVENTORY $ 7,70,200

=4,90,000+2,90,000-9,800

=7,70,200(add inventory values and less unreailsed profit)

SALES $ 15,82,000

=9,80,000+7,80,000-1,78,000

=15,82,000(add sales value and less intra entity value)

COST OF GOODS SOLD $ 7,11,800

=4,90,000+3,90,000-1,78,000+9,800

=7,11,800(add cost of goods sold values-intra entity transfer+unreailsed profit)

OPERATING EXPENSES $ 5,35,200

=2,25,000+2,95,000+,15,200

=5,35,200(add operating exp values and add amortization expense)

NET INCOME attribution to non controlling interest $ 14,000

=(7,80,000-3,90,000-2,95,000-15,200-9,800)*20%

=70,000*20/100

=14,000(20% of barone's net income , less amortization expense and less unrealised profit)

NOTE:

Customer list amortization=76,000/5yrs

=$15,200 per annum

Intra -entity gross profit = 1,78,000-1,29,000

=$49,000 (inventory remaining at year end)

Unrealised profit =49,000*20/100 (20% on unsold stock)

=$9,800